# Symphony-MoE: Harmonizing Disparate Pre-trained Models into a Coherent Mixture-of-Experts

### v1

@@ 5 Conclusion
This paper presents Symphony-MoE, a framework for building powerful Mixture-of-Experts (MoE) models by upcycling experts from multiple, diverse pre-trained sources. We tackle the core challenge of parameter space misalignment through a two-stage process: training-free functional alignment and lightweight router training. Experiments show that Symphony-MoE consistently outperforms baselines on multi-domain tasks and generalizes well to out-of-distribution data. This approach offers a scalable path to leverage the collective knowledge embedded in existing specialized models—beyond the limitations of single-model upcycling.
@@
Limitations and future work. Our current framework requires that all source models share an identical architecture. This constraint limits the pool of potential experts that can be integrated. Future work could focus on developing more advanced alignment techniques capable of harmonizing models with minor architectural dissimilarities, such as different activation functions or normalization strategies.
The quality of the functional alignment is contingent on the diversity and representativeness of the calibration dataset. While we proposed a simple and effective sampling strategy, optimizing the curation for a highly diverse set of experts is a non-trivial challenge.

### v2

@@ 5 Conclusion
This paper presents Symphony-MoE, a framework for building powerful MoE models by upcycling experts from multiple, diverse pre-trained sources. We tackle the core challenge of parameter space misalignment through a two-stage process. Experiments show that Symphony-MoE outperforms baselines on multi-domain tasks and generalizes well to out-of-distribution data. This approach offers a scalable path to leverage the collective knowledge embedded in existing specialized models.
@@
Limitations. Our current framework requires that all source models share an identical architecture and are all language models. This constraint limits the pool of potential experts that can be integrated. Future work could focus on developing more advanced alignment techniques capable of harmonizing models with minor architectural dissimilarities or modality differences. The effectiveness of functional alignment depends on the diversity of the calibration dataset. Although we propose a straightforward sampling strategy, optimizing the process to ensure a highly diverse set of experts presents a significant challenge.
Appendix A Computational Cost Analysis of Pairwise Inter-Expert Alignment
The scalability of the Symphony-MoE framework is critically dependent on the computational cost of its alignment stage. This section provides an analysis of the computational complexity of the Pairwise Inter-Expert Alignment process, demonstrating its efficiency and viability for large-scale models.
The alignment process consists of three main steps: 1) Activation Collection, 2) Permutation Matching, and 3) Weight Remapping. We analyze the cost of each step for a single layer, which can then be multiplied by the number of layers to determine the total cost.
1. Activation Collection: This step involves performing a forward pass through the anchor model ($M_{1}$) and each non-anchor model ($M_{i}$) to extract FFN activation matrices. The cost is proportional to the size of the calibration dataset ($D_{cal}$) and the number of parameters in the models. Let $C_{forward}$ be the cost of a single forward pass. For $N$ models, this step is performed for all N1N-1 non-anchor models. Thus, the total cost for this step is $(N-1)\times C_{forward}$. Since the forward pass is a standard operation in neural networks, this cost is manageable.
2. Permutation Matching: For each layer, we solve a linear assignment problem to find the optimal permutation matrix ($P_{i}^{(l)}$) that aligns the neurons of a non-anchor model to the anchor model. This is formulated as:
$\min_{P\in\mathcal{P}}\|A_{1}^{(l)}-A_{i}^{(l)}P\|_{F}^{2}$
where $\mathcal{P}$ is the set of permutation matrices. This problem is efficiently solved using the Hungarian algorithm. The computational complexity of the Hungarian algorithm for a $d\times d$ matrix, where $d$ is the number of neurons in the FFN layer, is $O(d^{3})$.
3. Weight Remapping: The computed permutation matrix is then applied to the weights of the FFN layers. This involves permuting the columns of the first linear layer’s weights ($W_{up,i}^{(l)}$) and the rows of the second linear layer’s weights ($W_{down,i}^{(l)}$).
$W_{up,i}^{\prime}=W_{up,i}^{(l)}P_{i}^{(l)}$
$W_{down,i}^{\prime}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$
The cost of these matrix multiplications for an FFN layer with input dimension $d_{in}$, hidden dimension $d$, and output dimension $d_{out}$ is $O(d_{in}\times d\times d)$ for the first layer and $O(d\times d\times d_{out})$ for the second. In typical transformer architectures, $d_{in}=d_{out}$, so the total cost for remapping is approximately $O(d^{3})$.
Overall Scalability: The dominant computational cost in the alignment of a single layer is the permutation matching and weight remapping, both of which are polynomial in the number of neurons ($d$). The total cost for aligning all FFN layers in N1N-1 non-anchor models is proportional to $(N-1)\times L\times O(d^{3})$, where $L$ is the number of layers. While the cubic complexity with respect to the number of neurons may seem significant, in practice, this alignment is a one-time, offline process performed before post-training. The cost is independent of the size of the training dataset for the final MoE model and is parallelizable across layers and models, making it a scalable solution for constructing large-scale MoE models from disparate pre-trained models.
Appendix B Experimental Setup Details
To create a set of specialized experts, we fine-tuned four distinct dense models. The foundation models include Qwen2.5-Base, its predecessor Qwen2-Base, and the code-specialized Qwen2.5-Coder. This selection was made to introduce disparity across model versions, pre-training data, and downstream tasks.
The instruction fine-tuning for all dense models was performed for 2 epochs with a consistent set of hyperparameters to ensure comparability.
For the model’s training configuration, we selected the AdamW optimizer with a learning rate set to 5e-5. During training, a batch size of 2 was used, and to prevent exploding gradients, we employed gradient clipping, limiting the max gradient norm to 1.0. The sequence cutoff length was set to 2048. Additionally, for regularization, a weight decay of 0.01 was applied, and a Cosine Decay scheduler was used to adjust the learning rate dynamically.
The calibration dataset $D_{cal}$, used in Stage 1, was constructed by sampling equally from each of the instruction fine-tuning datasets (See Table 5), resulting in a total of 10.404M tokens. This ensures a balanced representation of each expert’s domain during alignment and router training. For the post-training in Stage 2, we trained for 6 epochs using the AdamW optimizer with a learning rate of 5e-5.
Domain Datasets for Source Model Construction
General Alpaca
Math MetaMathQA
Code CodeAlpaca
Science SciQAG
Table 4: Explanation of instruction-finetuning datasets for source models. This batch of data was entirely used for training and was not sampled.
Domain Source of the Calibration Dataset
General SlimPajama-Wikipedia
Math Finemath
Code SlimPajama-Github
Science Scientific-Papers
Table 5: Explanation of the source of the Calibration Dataset $D_{cal}$. The number of tokens sampled from each data source is the same to ensure a balanced composition of $D_{cal}$. The random seed for sampling is set to 114514.
Appendix C Description and Reproduction Details of Baselines
MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
Dense (1B)
General ($M_{1}$, ANC) 61.35±0.39 35.14±1.09 45.05±0.55 40.15±3.80 30.11±1.60 42.36 31.55±0.20
Math ($M_{2}$) 58.10±0.39 38.95±1.33 47.05±0.54 39.50±3.79 32.05±1.62 43.13 31.89±0.20
Code ($M_{3}$) 45.20±0.41 37.10±1.33 38.22±0.55 46.88±3.81 27.85±1.51 39.05 29.81±0.19
Science ($M_{4}$) 57.90±0.40 34.20±1.30 37.95±0.53 38.14±3.74 33.05±1.57 40.25 30.98±0.19
MoE (1B $\times$ 4)
BTX 47.33±0.88 32.15±0.45 41.80±0.58 31.77±3.15 26.95±1.15 36.00 28.15±0.18
BAM 52.95±0.75 38.10±0.61 46.85±0.59 40.10±2.89 28.88±1.89 41.38 30.85±0.18
Drop 59.20±0.81 37.85±1.12 46.90±0.43 42.15±1.89 32.95±1.49 43.81 34.25±0.19
Symphony (Ours) 60.88±0.31 38.80±1.09 47.11±0.36 45.95±3.69 33.15±2.31 47.18 37.33±0.22
Table 6: Performance comparison of dense (Llama 3.2 1B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data.
All baselines in our comparison follow a common two-stage pipeline: upcycling → post-training. In the post-training stage, to ensure a fair comparison, all models are trained solely on the extended calibration dataset $D_{\mathrm{cal}}$ (as constructed in our framework), using the AdamW optimizer for 6 epochs with a learning rate of 5e-5. In the post-training phase, all modules are trainable, and the learning rate remains consistent.
The upcycling strategies adopted by each baseline are as follows:
(a) BTX constructs experts by directly reusing the FFN weights from each dense model. The shared backbone is formed by linearly averaging all model weights.
(b) BAM reuses FFN weights and partial attention weights ($W^{q}$, $W^{o}$) to construct experts, while the remaining weights are linearly averaged to form the shared backbone.
(c) Drop-Upcycling reuses FFN weights and applies Gaussian perturbation to randomly selected parameters to prevent expert homogenization. The remaining weights are reused and averaged to build the shared backbone, and the expert modules are also updated during training.
Appendix D Additional Experimental Results
For small-scale models, the Llama series uses Llama3.2 1B for all versions ($M_{1},M_{2},M_{3},M_{4}$) to isolate the impact of varying training tasks, data, and parameters on model performance.